OpenStudy (anonymous):

what is the integral of (x+3)/(x^2+1)

OpenStudy (anonymous):

Hint: break up the integral as x/(x^2 + 1) + 3/(x^2 + 1) The first integral is easy with u = x^2 + 1 Second integral involves arc tan u

OpenStudy (anonymous):

I got the first integral. I guess I am confused about the 3 in the second

OpenStudy (anonymous):

The second integral is 3 times the integral of 1/(x^2 + 1) = Arc tan (x), so the second integral is 3 Arc tan (x).

OpenStudy (anonymous):

The integral of dx/(1+x^2) = Arc tan (x)...thats one of them you should know

OpenStudy (anonymous):

The first integral is 1/2 ln(x^2 + 1)

OpenStudy (anonymous):

Okay that's what I got! Thanks so much

OpenStudy (anonymous):

Welcome.

OpenStudy (anonymous):

Your final answer should be: (1/2) ln (x^2 + 1) + 3 arc tan (x) + C